OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 2x + 4y = 8

OpenStudy (anonymous):

1st of all we can divide the whole equation by 2 - x + 2y = 4. 2y = -x + 4. y = -x/2 + 2 (dividing both sides of equation by 2) now can u find slope and y-intercept?

OpenStudy (anonymous):

thank u